OpenStudy (anonymous):

Which statements are true regarding the sequence below? Check all that apply. 10/3, 4 , 24/5, 144/25

OpenStudy (anonymous):

A) It can be represented using the formula f(x + 1) = 6/5(f(x)) when f(1)=10/3 B) It can be represented using the formula f(x)=4[6/5]x C) It can be represented using the formula f(x)= 10/3 [6/5]x-1 D) The domain of the sequence is all real numbers. E) The range of the sequence is all natural numbers.

OpenStudy (anonymous):

it is a geometric sequence find r try to find tn where n=x

OpenStudy (anonymous):

so its A,B,C,E?

OpenStudy (anonymous):

A is correct.

OpenStudy (anonymous):

so its A thats it?

OpenStudy (anonymous):

How B,C,E? show

OpenStudy (anonymous):

because it says to [ check all that apply]

OpenStudy (anonymous):

In C\[f(x)=\frac{ 10 }{ 3 }\left( \frac{ 6 }{ 5 } \right)^{x-1}\] but it is not exponential in your statement.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

if it is like this then it is also true

OpenStudy (anonymous):

in E 10/3,24/5,144/25 are not natural numbers.

OpenStudy (anonymous):

that will be real number then because those numbers are real

OpenStudy (anonymous):

it is range not domain

OpenStudy (anonymous):

as you have typed,only A is correct.

OpenStudy (anonymous):

A, C correct